# IrokoBench: A New Benchmark for African Languages in the Age of Large Language Models

### v1

@@ 3.4 LLMs used for evaluation
We evaluate on three encoder-decoder open LLM: Flan-T5-XXL [19], mT0-XXL-MT [42], and Aya-101 [59] that have been instruction fine-tuned on T5 (FlanT5) and multilingual T5 pre-trained on 101 languages (mT0 and Aya-101) models. Furthermore, these models are also all designed to be massively multilingual and explicitly optimized to work outside of English. The instruction data were constructed from several supervised NLP tasks with diverse prompts in different languages. Aya-101 additionally includes some synthetic data generated from Cohere Command model [20], and a large amount of translated data from English to cover several languages [55]. The languages covered during instruction tuning differ for different models, mT0, FlanT5, and Aya covered 46, 60, and 101 languages respectively.
Additionally we evaluate of seven decoder-only open LLM models: BLOOMZ 7B [65], Gemma 7B [56], LLaMa 2 7B [58], LLaMa 3 (8B & 70B) [41], Command-(R & R+) [20]. These models’ weights are openly available under various licenses, ranging from fully permissive to non-commercial, research-only licenses. We evaluate the instruction-tuned variant of these models. Generally, with the exception of BLOOMZ, these models did not report including any African languages in their pre-training. BLOOMZ included 22 African languages, mainly from Niger-Congo language family.
@@
We limit our evaluation to only OpenAI GPT (3.5-0125, 4-Turbo-0125, 4o) [46], and Claude OPUS [10] models. Recent work has shown that proprietary models tend to exhibit better multilingual capabilities [8], although specifics regarding their pre-training and instruction fine-tuning processes are not disclosed.
@@ 3.5 Evaluation settings
We conduct two types of evaluations: in-language evaluation and translate-test evaluation, where test instances are automatically translated into English using a machine translation engine. For automatic translation, we use NLLB-200 (3.3B) [57]. In both in-language and translate-test setups, we perform cross-lingual transfer experiments from English and zero-shot evaluations by prompting LLMs. Few-shot evaluations are performed only for the two best models (one open and one closed model) in the in-language setting.
We use the EleutherAI LM Evaluation Harness (lm-eval) tool [15]—a popular evaluation tool that is helping to standardize LLM evaluation, especially for open models on HuggingFace Model Hub. For closed models, except for the AfriMGSM task where we still use lm-eval, we employ a verbalizer [25, 52] for prediction and evaluation. We provide more details in Appendix A.2.
@@
We first conduct a study on cross-lingual transfer in a supervised learning setting by fine-tuning the English training data ($400K$ instances) from  Conneau et al. [21] and evaluating on the remaining languages. This experiment focuses solely on the NLI task due to the availability of training data for supervised learning. The evaluation employs several masked language models, including XLM-R [22], Serengeti [1], AfroXLMR-{base, large} [9], AfroXLMR-76L [4]. We report the result of the best model in the paper and others in Appendix A.5.
@@
In a zero-shot setting, we use the prompts detailed in Table 8. For few-shot evaluations, we conduct a 5-shot assessment for both AfriMMLU and AfriXNLI, and an 8-shot assessment for AfriMGSM.
Table 2: Main results: Average performance of various LLMs on AfriMMLU, AfriXNLI and AfriMGSM. Except for AfriMGSM, which uses the Exact Match metric, others use the Accuracy.
@@
AfroXLMR-76L 559M 65.7 63.6 65.7 63.6 88.2 83.3
FlanT5-XXL 11B 36.7 53.2 25.4 35.1 1.1 8.2 21.1 32.2 44.7 35.4
mT0-XXL-MT 13B 51.0 49.9 28.2 28.9 2.6 2.6 27.3 27.1 34.4 33.3
Aya-101 13B 49.9 49.4 30.3 31.1 3.3 6.6 27.9 29.0 36.2 35.0
BLOOMZ 7B 7B 39.4 47.5 23.5 26.4 4.1 2.1 22.4 25.3 32.0 29.9
Gemma 7B 7B 36.4 42.0 29.5 38.3 4.6 8.2 23.5 29.5 41.2 37.6
LLaMa 2 7B 7B 33.4 42.3 24.7 29.9 1.7 5.9 19.9 26.0 34.8 32.3
LLaMa 3 8B 8B 35.8 44.1 27.5 34.4 2.9 27.0 22.0 35.2 47.0 39.5
LLaMa 3 70B 70B 34.2 40.9 34.1 47.4 8.2 45.3 25.5 44.5 70.7 57.7
Command R 35B 34.0 46.1 27.0 36.5 1.6 5.8 20.9 29.5 42.1 38.4
Command R+ 104B 37.9 45.4 26.1 35.7 2.3 9.2 22.1 30.1 47.8 39.3
Claude OPUS UNK 58.1 56.4 43.0 47.6 25.9 32.7 42.3 45.6 73.3 64.0
GPT-3.5 Turbo UNK 41.5 52.2 31.8 39.6 5.3 19.3 26.2 37.0 64.7 54.1
GPT-4-Turbo UNK 60.0 52.8 50.9 47.0 26.8 25.1 45.9 41.6 81.7 74.1
GPT-4o UNK 62.7 59.9 52.3 52.3 29.3 32.4 48.1 48.2 72.5 65.1
Average 28.2 34.3 51.6 45.4
@@ 4 Results
Table 2 shows the result of zero-shot evaluation for various LLMs. On average, there is a significant performance gap between African languages and English (up to $45\%$) and French (up to $36\%$). The best model for African languages is GPT-4o, with an average performance of 48.1 across the evaluated tasks. However, GPT-4-Turbo outperforms GPT-4o for English and French, showing improvements of 9.2 and 9.0, respectively, across all tasks. It appears that GPT-4o’s enhancements for low-resource languages may negatively impact its performance on high-resource languages. Finally, as shown in Table 7, the languages with the lowest performance have the least monolingual data on the web.
@@
Our results, as presented in Table 2, indicate that the closed models Claude OPUS, GPT-4-Turbo, and GPT-4o consistently outperform the open models on our IrokoBench benchmarks. The closed models achieve average performance scores ranging from $42.3$ to $48.1$ across all tasks. The performance gap between the best closed model (GPT-4o) and the best open model (Aya-101) is $20.2$. Notably, the largest performance differences are observed in the AfriMMLU and AfriMGSM tasks, where GPT-4o outperforms LLaMa 3 70B by $18.2$ and $21.1$, respectively. For the AfriXNLI task, mT0-XXL-MT and Aya-101 perform better than LLaMa 3 70B and also achieve the highest scores among all open-weight models.
Majority models perform worse for in-language prompting Most users would prefer to prompt in their native language; however, we find that almost all models we benchmark perform better with prompts translated into English. Only a few exceptions, such as GPT-4-Turbo, mT0-XXL-MT, and Aya-101, perform better in the in-language evaluation. Among open-source models, Aya-101 performs the best in the in-language evaluation, surpassing much larger open-weight models like LLaMa 3 70B. Specifically, LLaMa 3 70B and LLaMa 3 8B benefit the most from the translate-test approach, showing average improvements of $19.0$ and $13.1$, respectively. Notably, LLaMa 3 70B achieves the best overall results for the AfriMGSM task with the translate-test, outperforming GPT-4o by more than $12.9$. Similarly, we observe significant performance gains for the FlanT5-XXL and GPT-3.5 Turbo models on the AfriXNLI and AfriMGSM tasks when using the translate-test. We attribute this boost to the fact that these models are heavily English centric.
Table 3: AfriMMLU results in in-language and translate-test scenarios: Option prediction accuracy per language. Average computed on only African languages.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
@@
Aya 37.2 32.0 30.4 30.8 31.6 29.2 32.6 29.2 29.0 33.4 33.8 31.6 30.8 25.4 32.8 32.6 32.4 31.1
LLaMa 3 8B 43.4 38.4 35.2 34.4 33.2 36.0 36.0 27.8 34.2 31.4 38.8 43.6 29.0 29.6 36.6 33.2 33.6 34.4
LLaMa 3 70B 64.4 54.4 42.2 44.4 48.6 47.2 49.8 41.2 49.6 48.4 54.0 60.0 41.4 29.8 49.6 49.2 49.2 47.4
GPT-4 Turbo 61.0 50.0 37.6 43.2 47.6 48.2 46.8 40.4 44.6 46.2 53.0 72.7 36.4 28.0 48.4 46.6 49.0 47.0
GPT-4o 67.2 57.6 40.6 49.4 56.4 55.0 50.2 41.8 51.6 49.8 59.0 80.0 37.2 33.0 53.0 52.4 54.2 52.3
Table 4: AfriMGSM results in in-language and translate-test scenarios: Exact Match score per language. Average computed on only African languages
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
Prompt LLMs in African Language
Aya 24.8 22.0 7.6 2.8 5.6 4.0 6.8 2.0 4.4 2.4 3.2 3.2 11.2 1.2 2.4 2.0 3.6 4.0 4.2
LLaMa 3 8B 24.8 22.0 7.6 2.8 4.8 4.4 5.6 2.0 4.8 2.4 2.4 3.2 11.2 1.2 2.4 1.6 3.2 3.6 4.0
LLaMa 3 70B 86.4 54.4 5.2 2.8 15.6 9.2 8.4 5.6 8.0 4.0 4.8 6.0 34.8 3.2 4.0 4.4 7.2 7.6 8.2
GPT-4 Turbo 79.6 64.0 8.4 6.8 40.8 21.6 41.6 30.8 24.4 31.2 38.4 38.0 54.4 6.8 9.2 31.2 29.2 31.2 27.8
GPT-4o 58.4 43.6 4.4 5.6 42.0 32.0 36.4 30.8 22.8 39.2 33.2 28.8 44.8 10.4 12.4 29.2 30.8 28.4 27.0
Translate-Test (Eval. in English)
Aya 8.0 8.0 6.0 7.6 4.0 7.2 5.2 4.4 5.2 8.4 8.0 10.8 4.4 5.6 6.0 8.4 6.4 6.6
LLaMa 3 8B 44.0 31.6 18.0 27.2 24.0 34.0 27.2 23.2 28.8 26.8 33.6 38.0 19.6 12.8 28.8 29.6 28.8 27.0
LLaMa 3 70B 69.6 53.6 27.6 48.4 40.4 54.0 47.6 34.0 47.6 46.8 54.0 69.6 29.2 20.8 50.0 49.2 51.2 45.3
GPT-4 Turbo 63.6 42.0 20.4 39.2 35.2 47.6 41.6 27.2 38.8 39.2 46.0 56.8 23.2 12.4 41.2 42.0 40.4 37.1
GPT-4o 52.0 18.4 37.6 36.4 28.0 34.8 38.4 28.4 32.8 32.4 36.8 45.2 17.6 12.4 31.6 36.0 38.0 31.6
Table 5: AfriXNLI results in in-language and translate-test scenarios: Option prediction accuracy per language. Average computed on only African languages.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
@@
Aya 60.3 55.2 47.0 51.8 52.2 50.3 33.0 49.0 49.0 52.0 53.3 54.0 45.5 41.5 54.5 48.8 54.0 49.4
LLaMa 3 8B 52.2 48.3 43.3 43.2 44.8 45.3 33.7 46.0 45.2 45.5 47.5 47.5 40.5 38.7 44.5 44.2 47.0 44.1
LLaMa 3 70B 48.5 42.8 41.3 43.7 39.8 40.3 33.3 43.3 39.7 41.1 43.0 44.3 39.7 38.7 42.3 38.2 42.7 40.9
GPT-4 Turbo 75.2 64.0 43.0 52.8 59.0 55.8 31.7 51.0 52.5 55.8 58.5 60.2 45.0 39.7 63.2 52.5 59.5 52.8
GPT-4o 80.0 69.0 50.7 62.2 63.3 61.0 33.0 63.2 58.7 65.8 67.0 69.5 51.3 46.2 69.7 60.5 67.3 59.9
Table 6: Ablation results: Effect of using different prompts for AfriXNLI. Best results per prompt are in bold. Average computed on only African languages.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul avg
ANLI prompt [44]—simple prompt
Aya 61.5 60.5 58.2 42.2 56.8 54.7 48.7 34.5 44.5 51.7 55.8 54.7 55.2 48.2 37.3 53.7 48.3 54.3 49.9
Gemma 7B 51.5 47.8 34.3 35.3 38.8 39 35.8 33.2 34.8 35.7 38.8 35.7 39.3 37.2 35.8 37.2 36.5 35.7 36.4
LLaMA 3 70B 51.3 50.2 36.8 33.3 35.3 36.5 34.5 33 33.5 33.3 33.8 34.3 36.5 33 33.5 34 33 33 34.2
Opus 69.3 67.5 54.8 41.3 55.5 52.8 46.8 27.8 51.3 52.7 50.2 56.8 57.8 46.7 41.0 57.7 53.3 54.5 50.1
GPT-4o 79.8 74.5 61.8 42.0 61.0 58.2 55.8 33.2 58.0 58.3 62.2 62.0 65.5 49.2 46.5 62.5 55.2 59.8 55.7
Lai’s prompt [37]—detailed prompt
Aya 49.4 45.1 47.1 32.3 44.2 43.8 39.1 21.5 40.3 38.5 41.7 43.7 43.7 38 27.3 46.2 39.8 43.2 39.4
Gemma 7B 64 58.2 19.5 19.1 24.2 25.8 22.2 24.2 20.2 20.6 25.8 20.3 46.7 19.6 19 31.1 20.3 25.8 24.0
LLaMA 3 70B 72.8 51.4 41.9 27.5 47.1 43 39.8 23.9 38 38.2 37.9 37.4 52.1 22.1 29.1 39 36.7 36.3 36.9
Opus 85.7 74.7 61.3 54.5 61.5 45.0 64.5 30.5 63.7 50.2 57.0 68.3 70.5 56.0 50.0 68.8 63.7 63.7 58.1
4.1 Task-specific results
Here, we examine the individual language performance per task, comparing which task prefers in-language v.s. translate-test evaluation. We compare the performance on a subset of LLMs: Aya-101, LLaMa 3 (8B & 70B), GPT-4-Turbo, and GPT-4o. Other LLMs are in Appendix A.4.
AfriMMLU evaluation is better when prompting in-language
Table 3 shows the result of different LLMs on AfriMMLU using in-language vs. translate-test. We found in-language prompting to be better on 10 out of the 16 African languages we evaluated; ibo, orm, wol, and yor have comparable performance with two approaches. However, we find a large improvement in the performance of the translate-test for swa and ewe with a boost of $12.4$ (GPT-4o comparison) and $10.4$ (LLaMa 3 70B comparison), respectively. For Swahili (swa), this improvement may be due to the strong machine translation performance on NLLB-200 , as reported in Adelani et al. [2] and Wang et al. [62].
AfriMGSM performance receives significant boost with translate-test evaluation
AfriMGSM is our most difficult benchmark in terms of overall performance on in-language evaluation. Table 4 shows that we can achieve a significant boost in performance with the translate-test on all languages we evaluated. We hypothesize that the current LLMs are better at reasoning in English than any other language. Interestingly, LLaMa 3 70B has the overall best performance boost of approximately $82\%$, and the performance boost confirms our hypothesis since it is more English-centric.
Cross-lingual transfer in-language achieves better results
When there is large enough labeled data in English, we could leverage this cross-lingual signal for zero-shot evaluation. We trained on 400k English NLI examples, and we performed zero-shot transfer in in-language and translate-test setting. Table 5 shows that cross-lingual transfer using an Africa-centric smaller language model (AfroXLMR-76L) gave better results than prompting LLMs on average. AfroXLMR-76L has been pre-trained on all languages in IrokoBench, which explains the impressive performance. However, for multilingual encoders that have not seen some of the languages, prompting GPT-4o seems to be better, as shown in Appendix A.5.
Impact of few-shot vs zero-shot
Figure 2 shows the few-shot results for the IrokoBench datasets leveraging Aya-101, LLaMa 3 70B and GPT-4o when we provide few examples in in-language setting. We found out that Aya-101 LLM consistently benefited the least from additional few shots examples in all cases with only $+1.4$, $+2.0$ and $+0.1$ respectively. On the other hand, we see a large boost LLaMa 3 70B on both AfriXNLI and AfriMMLU tasks with $12.4$ and $6.5$ points improvement respectively. We hypothesize that additional examples in the language LLaMa 3 70B was not trained on were particularly helpful for the classification tasks. However, for mathematical reasoning tasks, this was unhelpful since LLaMa 3 70B is only able to reason properly in English as shown in Table 4. In general, GPT-4o benefits from adddtional few-shot examples, there is large improvement for AfriMGSM ($+10.5$) but smaller points improvement on AfriXNLI and AfriMMLU are smaller with $+2.0$ and $+0.4$ respectively.
One possibility for the difference in added performance from few-shot in part is due to the difference in context length. Aya-101 was by necessity built upon the mT5 [66] pre-trained base model given it was one of the few pre-trained models that had been trained on 101 languages. mT5 is relatively outdated given the rapid advances in LLM technology since its release in 2019. In particular, the context length of Aya-101 is only 1024 versus 8000 for LLaMa 3 70B and 32000 for GPT-4o. Hence, the truncation of some few-shot examples may disproportionately impact Aya-101.
Sensitivity to prompt templates
To understand whether sensitivity to prompts impacts results we perform an ablation for NLI where we evaluate results given two different prompts. Table 6 shows the results of two prompts we tested for: the first one is simpler [44], where the model is required to output true, false or neither, while the second prompt provides a detailed task description [37], and the output we are expecting is the label entailment. contradiction or neutral. We found out that open models like Aya-101 and Gemma 7B prefer the simpler prompt with Aya-101 achieving close to $50.0$ accuracy points on the task on average. However, closed models like Claude OPUS and GPT-4o prefer the Lai et al. [37] detailed prompt with $8.0$ and $7.0$ improvement in performance over the simpler prompt. In general, regardless of the prompt selection, GPT-4o is better than open models, but we can achieve better results for open models by providing a more detailed task description. While this evaluation is only for two prompts, an extensive prompt search would be important to explore all the tasks we evaluated.

### v2

@@ 3.4 LLMs used for evaluation
We evaluate on two encoder-decoder open LLM: mT0-XXL-MT (Muennighoff et al., 2023), and Aya-101 (Ustun et al., 2024) that have been instruction fine-tuned and multilingual T5 pre-trained on 101 languages (mT0 and Aya-101) models. Furthermore, these models are also all designed to be massively multilingual and explicitly optimized to work outside of English. The languages covered during instruction tuning differ for different models, mT0 and Aya covered 46 and 101 languages respectively.
Additionally we evaluate on eight decoder-only open LLM models: BLOOMZ 7B (BigScience-Workshop et al., 2023), Gemma 2 (9B & 27B) Riviere et al. (2024), LLaMa 3 8B (Meta, 2024), LLaMa 3.1 (8B & 70B) (Dubey et al., 2024), Command-R (August version) (Cohere, 2024), and LLaMaX 3 8B (Lu et al., 2024). These models’ weights are openly available under various licenses, ranging from fully permissive to non-commercial, research-only licenses. We evaluate the instruction-tuned variant of these models. All models were pre-trained from scratch except LLaMaX that undergo continue pre-training on 100 languages including 13 languages covered in IrokoBench, except Ewe, Twi, Kinyarwanda and Vai. We used LLaMAX3-8B-Alpaca, instruction-tuned on English Alpaca Taori et al. (2023).
@@
We limit our evaluation to only OpenAI GPT (3.5-0125, 4-Turbo-0125, 4o-mini-07-18, 4o-08-06) (OpenAI, 2024), Gemini-1.5-Pro (Reid et al., 2024), and Claude OPUS (Anthropic, 2024) models. Recent work has shown that proprietary models tend to exhibit better multilingual capabilities Ahuja et al. (2023b), although specifics regarding their pre-training and instruction fine-tuning processes are not disclosed.
@@ 3.5 Evaluation Settings
We conduct two types of evaluations: in-language and translate-test evaluation, where test instances are automatically translated into English using a machine translation (MT) engine. For MT, we use NLLB-200 (3.3B) (Costa-jussà et al., 2024). In both in-language and translate-test setups, we perform cross-lingual transfer experiments from English and zero-shot evaluations by prompting LLMs. Few-shot evaluations are performed only for the three best models (two open and one closed) in the in-language setting. For AfriMGSM in both settings, we use Chain-of-Thought (COT) reasoning.
We use the EleutherAI LM Evaluation Harness (lm-eval) tool (Biderman et al., 2024)—a popular evaluation tool that is helping to standardize LLM evaluation, especially for open models on HuggingFace Model Hub. For closed models, we employ a verbalizer Gao et al. (2021); Schick and Schütze (2021) for prediction and evaluation. All models are prompted with five different templates. We provide more details in Appendix A.2.  [^3] We make use of Cohere API for Command-R inference.
@@
We first conduct a study on cross-lingual transfer in a supervised learning setting by fine-tuning the English training data ($400K$ instances) from  Conneau et al. (2018) and evaluating on the remaining languages. This experiment focuses solely on the NLI task due to the availability of training data for supervised learning. The evaluation employs several masked language models, including XLM-R Conneau et al. (2020), Serengeti Adebara et al. (2023), AfroXLMR-{base, large} Alabi et al. (2022), AfroXLMR-76L Adelani et al. (2024). We report the result of the best model in the paper and others in Appendix A.6.
@@
In a zero-shot setting, we use the prompts detailed in subsection A.2. For few-shot evaluations, we conduct a 5-shot assessment for both AfriMMLU and AfriXNLI, and an 8-shot assessment for AfriMGSM.
@@
AfroXLMR-76L 559M 65.7 63.6
mT0-XXL-MT 13B 51.0 49.9 27.9 28.4 2.9 2.5 27.3 26.9 34.5 33.1
Aya-101 13B 51.5 50.2 29.7 31.1 4.6 7.9 28.6 29.7 39.3 35.3
BLOOMZ 7B 7B 39.4 47.6 24.1 27.9 1.7 1.9 21.7 25.8 31.4 28.9
LLaMa 3 8B 8B 35.4 38.2 28.1 31.8 3.9 33.4 22.5 34.4 52.8 45.8
LLaMa 3.1 8B 8B 36.6 43.6 31.1 41.1 7.2 30.1 24.9 38.2 57.5 53.4
LLaMaX 3 8B 8B 40.8 33.3 29.3 35.2 4.8 9.2 24.9 25.9 38.5 33.1
Gemma 2 9B 9B 40.3 43.3 35.4 44.7 19.8 39.4 31.9 42.5 64.6 58.2
Gemma 2 27B 27B 42.8 49.0 39.9 48.8 28.5 46.1 37.1 48.0 76.3 69.9
LLaMa 3.1 70B 70B 38.0 42.8 39.4 51.3 24.6 45.6 34.0 46.5 73.5 65.5
Command-R 35B 43.4 57.0 27.8 40.8 5.7 38.3 25.6 45.4 71.7 63.0
Claude Opus UNK 58.1 56.4 43.0 47.6 25.3 32.7 42.3 45.6 73.3 64.0
Gemini-1.5-Pro UNK 59.4 49.9 60.2 53.1 55.4 44.3 58.3 49.1 82.6 75.3
GPT-3.5-Turbo UNK 42.1 45.5 38.1 46.8 10.6 37.1 30.2 43.1 71.9 62.9
GPT-4o-mini UNK 54.2 56.7 45.5 50.2 35.4 42.3 45.0 49.7 84.7 76.7
GPT-4-Turbo UNK 59.5 57.0 54.2 52.1 45.2 43.5 52.9 50.9 84.6 76.5
GPT-4o UNK 64.3 52.1 60.0 54.1 52.6 42.6 59.0 49.6 86.9 78.1
33.9 40.2 62.8 56.3
Table 2: Main results: Average performance of various LLMs on all tasks (ave. excl. eng, fra, and vai). Except for AfriMGSM, which uses the Exact Match metric, others use the Accuracy. The best result is in Bold and second best underlined. The top-2 open and closed models are in gray. We report only the best prompt (others in appendix).
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul ave
Prompt LLMs in African Language
Aya-101 (t2) 40.0 36.6 31.6 25.4 33.4 36.8 30.8 27.8 28.0 26.2 28.2 31.8 32.2 26.8 25.2 32.0 28.4 29.8 29.7
Gemma 2 27B (t1) 75.6 66.4 40.6 32.4 43.2 44.2 40.2 38.2 32.6 33.6 44.6 41.8 56.0 35.6 30.4 42.0 41.0 42.2 39.9
LLaMa 3.1 70B (t1) 76.4 69.4 41.6 32.2 47.6 47.2 38.6 40.0 34.4 35.6 41.6 39.0 55.8 28.4 31.6 34.2 41.4 40.4 39.4
GPT-4o (t1) 87.4 83.2 59.8 33.6 67.2 67.2 64.2 61.0 52.8 61.0 67.6 67.4 77.4 43.2 37.8 70.2 61.2 68.2 60.0
Translate-Test (Eval. in English)
Aya-101 (t2) 37.8 32.4 28.6 31.0 31.6 31.8 33.6 27.2 28.6 32.4 34.2 31.2 31.8 26.6 30.8 34.2 32.2 31.1
Gemma 2 27B (t1) 62.4 57.6 40.4 50.0 50.0 50.6 47.0 42.8 46.6 49.8 55.6 59.8 39.8 31.2 55.2 52.6 51.6 48.8
LLaMa 3.1 70B (t1) 67.4 55.6 44.8 50.6 55.8 55.6 53.8 46.8 49.4 53.6 59.6 63.0 41.2 32.8 55.4 49.6 52.8 51.3
GPT-4o (t1) 76.4 62.8 43.8 54.0 57.4 58.2 54.4 46.4 54.8 55.6 63.2 67.4 44.0 32.0 62.2 52.4 57.4 54.1
Table 3: AfriMMLU results in in-language and translate-test scenarios: Option prediction accuracy per language. Average computed on only African languages. The best prompt template for each model is in bracket.
@@ 4.1 Overall Results
Table 2 shows the result of zero-shot evaluation for various LLMs. On average, there is a significant performance gap between African languages and English (up to $28\%$) and French (up to $19\%$) on the best LLM. The best LLM for African languages is GPT-4o, with an average performance of 59.0 across the evaluated tasks. Finally, as shown in Table 9, the languages with the lowest performance have the least monolingual data on the web.
@@
Our results, as presented in Table 2, indicate that the closed models Claude Opus, Gemini-1.5-Pro, GPT-4-Turbo, and GPT-4o consistently outperform the open models on IrokoBench. The top-2 closed models achieve average performance scores ranging from $52.9$ to $59.0$ across all tasks. The performance gap between the best closed model (GPT-4o) and the best open model (Gemma 2 27B) is $21.9$. Notably, the largest performance differences for in-language setting are observed in the AfriMMLU and AfriMGSM tasks, where GPT-4o outperforms Gemma 2 27B by $20.1$ and $24.1$, respectively. For the AfriXNLI task, mT0-XXL-MT and Aya-101 perform better than bigger models like Gemma 2 27B and LLaMa 3.1 70B with 27B and 70B parameters respectively.
Majority of models perform worse for in-language prompting
Most users would prefer to prompt in their native language; however, we find that almost all models we benchmark perform better with prompts translated into English. Only a few exceptions, such as GPT-4-Turbo, and GPT-4o, perform better in the in-language evaluation. Specifically, Command R and LLaMa 3.1 8B benefit the most from the translate-test approach, showing average improvements of $+19.8$ and $+13.3$, respectively. Notably, Gemma 2 27B achieves the best overall results for the AfriMGSM task with the translate-test, outperforming GPT-4o by $+3.5$. We attribute this boost to the fact that these models are heavily English-centric.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul ave
Elementary Mathematics 88.0 83.0 76.0 55.0 73.0 84.0 77.0 71.0 80.0 81.0 75.0 78.0 87.0 68.0 61.0 76.0 80.0 80.0 75.1
Global Facts 67.0 60.0 44.0 32.0 56.0 49.0 60.0 54.0 41.0 52.0 48.0 49.0 59.0 33.0 36.0 56.0 42.0 57.0 48.0
High School Geography 93.0 86.0 56.0 25.0 67.0 64.0 57.0 55.0 48.0 54.0 69.0 68.0 76.0 34.0 27.0 63.0 53.0 73.0 55.6
High School Microeconomics 98.0 90.0 54.0 26.0 70.0 58.0 55.0 63.0 44.0 55.0 59.0 62.0 82.0 32.0 19.0 78.0 51.0 61.0 54.3
International Law 90.0 91.0 64.0 36.0 71.0 75.0 75.0 76.0 55.0 65.0 78.0 72.0 86.0 45.0 38.0 84.0 75.0 76.0 66.9
Average 87.2 82.0 58.8 34.8 67.4 66.0 64.8 63.8 53.6 61.4 65.8 65.8 78.0 42.4 36.2 71.4 60.2 69.4 60.0
Table 4: GPT-4o AfriMMLU results by subjects: Option prediction accuracy per language. Languages with at least 60% accuracy in four subjects are in Cyan.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul ave
@@
Aya-101 (t1) 8.4 8.4 6.4 7.6 6.0 10.0 6.4 6.8 7.6 6.8 10.4 9.2 6.0 8.4 8.8 8.8 8.0 7.9
Gemma 2 27B (t4) 70.8 53.2 30.0 54.0 44.0 55.2 47.2 34.4 46.0 48.0 54.4 69.6 29.2 21.6 48.4 48.4 54.0 46.1
LLaMa 3.1 70B (t4) 73.6 54.8 30.4 48.0 43.2 52.8 48.0 35.6 44.0 46.8 55.2 72.0 26.4 20.4 49.6 48.4 54.0 45.6
GPT-4o (t2) 70.0 48.4 23.6 46.8 39.2 51.6 44.8 35.2 42.0 42.4 54.8 68.0 26.4 16.0 45.6 47.2 49.2 42.6
Table 5: AfriMGSM results in in-language and translate-test scenarios: Exact Match score per language. Average computed on only African languages. The best prompt template for each model is in bracket.
Model eng fra amh ewe hau ibo kin lin lug orm sna sot swa twi wol xho yor zul ave
@@
Aya-101 (t4) 61.2 60.7 40.8 53.8 52.3 50.3 33.0 48.7 51.3 54.0 56.0 54.0 44.5 39.3 56.8 51.0 57.0 50.2
Gemma 2 27B (t2) 60.8 55.8 45.5 48.5 49.7 47.7 31.5 51.3 52.3 51.7 50.2 55.7 47.2 40.2 54.8 50.2 51.5 49.0
LLaMa 3.1 70B (t2) 58.8 42.8 39.3 45.2 39.8 36.2 32.7 44.3 46.0 51.8 50.8 52.5 38.8 34.5 44.8 41.8 42.8 42.8
GPT-4o (t3) 73.8 63.8 42.8 54.8 53.3 52.8 32.3 54.3 55.5 58.2 58.0 58.5 45.7 38.8 58.2 53.5 52.8 52.1
Table 6: AfriXNLI results in in-language and translate-test scenarios: Option prediction accuracy per language. Average computed on only African languages. The best prompt template for each model is in bracket.